a duopoly market in which the demand is = p = 120 - 2Q. The total cost for firm 1 is TC1=20q1, the total cost for firm 2 is TC1=40q2. The good is homogeneous. The two firms collude. What is the equilibrium price?
Answer #1

A duo poly market in which the demand ist _P=120-20 where Q = 91 +2 = 120- 2191192) Total revenue (TR) = P(qi+q2) =1120-2 (2,- -227-29 - 4 q qz +100g, +80.22 Now; _2732421 – 4227100=0:1 22 - (1) -4844422 - 1oo And 271 . -422-42,790 0 222 - E S ) Way